OpenStudy (anonymous):

what is the circumference of a circle with diameter of 7cm?

OpenStudy (anonymous):

u know it

OpenStudy (anonymous):

circumference=2 pie r radius=diameter/2

OpenStudy (anonymous):

so 21.98?

OpenStudy (anonymous):

\[7\pi =21.99 \text{ cm} \]

OpenStudy (anonymous):

its not showing what you put

OpenStudy (anonymous):

The ratio of the a circle's circumference to it's diameter is Pi.\[\frac{c}{d}= \pi , \pi d = c \]